# Tora: Trajectory-oriented Diffusion Transformer for Video Generation

### v3

@@ Introduction
VideoComposer (Wang et al. 2023) 529 668 856 0.2335 0.2284 0.2236 15.11 29.14 58.76
DragNUWA (Yin et al. 2023) 475 593 784 0.2385 0.2341 0.2305 10.04 17.33 41.25
AnimateAnything (Dai et al. 2023) 487 602 775 0.2399 0.2342 0.2313 13.39 27.28 51.33
TrailBlazer (Wan-Duo Kurt Ma 2023) 459 581 756 0.2403 0.2351 0.2322 11.68 19.47 44.10
MotionCtrl (Wang et al. 2024b) 463 572 731 0.2412 0.2376 0.2331 9.42 16.46 38.39
@@
Table 1: Quantitative comparisons with state-of-the-art motion-controllable video generation models. As the number of generated frames increases, Tora demonstrates a growing performance advantage over the UNet-based methods, maintaining a high degree of stability in trajectory control.
Data Processing and Training Strategy
Data Processing. We employ a structured data processing method to obtain high-quality training videos with consistent object motion. Initially, raw videos are segmented into shorter clips based on scene detection [^2] https://github.com/Breakthrough/PySceneDetect. Subsequently, we remove invalid videos, such as those with encoding errors, a duration of zero, and low resolution. Furthermore, we utilize aesthetic [^3] https://github.com/christophschuhmann/improved-aesthetic-predictor and optical flow scores (Xu et al. 2023) to filter out low-quality videos. To concentrate on the motion of primary objects, we implement camera motion filtering, using results from motion segmentation (Zhao et al. 2022) and camera detection to exclude instances predominantly exhibiting camera movement. Dramatic object motions in certain videos can lead to significant optical flow deviations, which may interfere with trajectory training. To address this, we retain these videos based on a probability of $(1-score/100)$. For eligible videos, we generate captions using the PLLaVA model (Xu et al. 2024). During inference, we utilize GPT-4o (OpenAI 2023) to refine prompts, ensuring alignment with training process for consistency. More detailed information about the filtering and the refinement process can be found in the supplementary materials.
Motion condition training. Inspired by DragNUWA and MotionCtrl, we adopt a two-stage training approach for trajectory learning. In the first stage, we extract dense optical flow (Xu et al. 2023) from the training video as the trajectory, providing richer information to enhance motion learning. In the second stage, we adjust the model from complete optical flow to more user-friendly trajectories by randomly selecting 1 to $N$ object trajectories based on motion segmentation results and flow scores. To improve the scattered nature of sparse trajectories, we apply a Gaussian filter for refinement. After completing the two-stage training, Tora facilitates flexible motion control using arbitrary trajectories.
Image condition training. We follow the mask strategy employed by OpenSora to support visual conditioning. Specifically, we randomly unmask frames during training, and the video patches of the unmasked frames are not subjected to any noise. This enables our Tora model to seamlessly integrate text, images, and trajectories into a unified model.
Experiments
Experimental Setup
Implementation Details. Tora is initialized with OpenSora v1.2 weights, and training videos have resolutions from 144p to 720p and frame counts ranging from 51 to 204. To balance training FLOP and memory usage, we adjust the batch size from 1 to 50. We use Adam Optimizer (Kingma and Ba 2015) with a learning rate of $2\times 10^{-5}$ on 4 NVIDIA A100. The 3D VAE in TE is initially trained on datasets  (Mehl et al. 2023; Mayer et al. 2016; Ranjan et al. 2020; Cabon, Murray, and Humenberger 2020) annotated with optical flow and then frozen during Tora training. We train Tora for 2 epochs with dense optical flow and fine-tune for 1 epoch with sparse trajectories. The maximum number of sampling trajectories $N$ is set to 16. The inference step and the guidance scale are set to 30 and 7.0, respectively.
Dataset. Our training videos are sourced from three datasets: 1) Panda-70M (Chen et al. 2024), from which we use the training-10M subset containing high-quality videos; 2) Mixkit (Envato 2024); and 3) Internal videos. The internal videos are manually annotated, with each clip labeled to include object masks and camera movement. Following our data processing pipeline, we select about 630k eligible videos for the training dataset. For inference, we curate 185 clips with diverse motion trajectories and scenes, to serve as a new benchmark for evaluating the motion controllability.
Metrics. We leverage standard metrics such as Fréchet Video Distance (FVD) (Unterthiner et al. 2018), and CLIP Similarity (CLIPSIM) (Wu et al. 2021) to quantitatively evaluate video quality. For assessing motion controllability, we leverage Trajectory Error (TrajError), which computes the average distance between the generated and pre-defined trajectories. Human evaluation is also introduced in supplementary materials due to space limitations.
Results
We compare our method with popular motion-guided video generation approaches in three settings: 16, 64, and 128 frames, all at a resolution of 512 × 512 for a fair evaluation. The provided trajectories are adjusted to fit the different video lengths. For most U-Net-based methods, we use sequenced inference, where the last frame generated from one batch serves as the visual condition for the next, aligning with their inference strategies. As shown in Table 1, in the 16-frame setting typical for U-Net methods, MotionCtrl and DragNUWA align better with the provided trajectories but still fall short compared to our proposed Tora. With an increasing number of frames, U-Net methods exhibit significant deviations, leading to misalignment errors that result in deformations, motion blur, or object disappearance in later sequences. In contrast, Tora remains highly robust across varying frame counts due to the transformer’s scaling abilities. In the 128-frame test setting, Tora’s trajectory accuracy surpasses other methods by 3 to 5 times, showcasing its outstanding motion control capabilities. Figure 4 presents an analysis of Trajectory Error across different resolutions and durations. Unlike U-Net models, which exhibit substantial trajectory errors over time, Tora shows only a gradual increase in error as duration extends. This aligns with the decrease in video quality observed in the DiT model. The results demonstrate that our method effectively maintains trajectory control over longer durations.
@@ Results
Figure 4: Comparison of Trajectory Error across various resolutions and durations. Unlike U-Net models, our method maintains motion control with a gradual increase in error.
Ablation study
We conduct several ablation studies to analyze the effects of our design choices. All models are evaluated using 480p resolution, a 16:9 aspect ratio, and 204 frames.
Trajectory Compression. To integrate the trajectory vector into the same latent space as video patches, we explore three different methods for trajectory compression, as summarized in Table 2. The first method samples the mid-frame as a keyframe for successive 4-frame intervals and uses patch-unshuffle for spatial compression. While simple, this approach is sub-optimal for motion control due to potential flow estimation errors during rapid movements or occlusions, and the increased dissimilarity between patches complicates learning. The second method employs average pooling to gather information from successive frames. Although this captures a general sense of movement, it sacrifices precision by averaging the trajectory’s direction and magnitude, diluting important motion details. To better preserve trajectory information between consecutive frames, we utilize a 3D VAE to extract the global context of successive trajectory intervals. The trajectory data is converted into RGB images to leverage existing 3D VAE weights. Extensive training on a large dataset of trajectory videos with this method yields the best results, highlighting the effectiveness of our customized 3D VAE approach for trajectory compression.
@@
Table 2: Evaluations of the impact of different trajectory compression methods.
Block design and integrated position of MGF. We train the three variant MGF blocks as previously described, with the results presented in Table 3. Notably, the adaptive norm block achieves lowest FVD and Trajectory Error, while also exhibiting the highest computational efficiency. This advantage stems from its ability to dynamically adapt features based on varying conditions without needing strict alignment, a common challenge with cross-attention. Additionally, it maintains temporal consistency by modulating conditional information over time, which is essential for incorporating motion cues. In contrast, channel concatenation can lead to information congestion, making motion signals less effective. We find that initializing the normalization layer as the identity function is vital for optimal performance. Additionally, placing the MGF module within the Temporal DiT block significantly enhances trajectory motion control, evidenced by a drop in Trajectory Error from 23.39 to 14.25.
@@
Training Strategies. We evaluate the two-stage training approach, with results in Table 4. Training only with dense optical flow is ineffective, as it fails to capture the details of sparse trajectories. Conversely, using only sparse trajectories provides limited information, complicating the learning process. By first training with dense flows and then fine-tuning with sparse trajectories, our model demonstrates better adaptability and versatility in managing various motion patterns, leading to improved overall performance.
@@
Table 4: Ablation of the type of training trajectories. “Hybrid” denotes the two-stage training strategy.
Figure 5: Qualitative Comparisons on Trajectory Control. All methods are capable of generating objects that follow the given trajectory. However, Tora not only adheres precisely to the specified trajectory but also produces smoother movement that conforms to the physics world.
Conclusion
This paper introduces Tora, the first trajectory-oriented Diffusion Transformer framework for video generation. Tora effectively encodes arbitrary trajectories into spacetime motion patches, which align with the scaling properties of DiT, thereby enabling more realistic simulations of physical world movements. By employing a two-stage training process, Tora achieves motion-controllable video generation across a wide range of durations, aspect ratios, and resolutions. Remarkably, it can generate high-quality videos that adhere to specified trajectories, producing up to 204 frames at 720p resolution. This capability underscores Tora’s versatility and robustness in handling diverse motion patterns while maintaining high visual fidelity. We hope our work provides a strong baseline for future research in motion-guided Diffusion Transformer methods.
@@ Conclusion
Qualitative Comparisons
While the main text focuses on quantitative comparisons with the motion-controllable video generation models and ablation studies on different designs for TE and MGF, here we provide further visual comparisons.
Compare with motion-controllable methods
@@
Figure 5 provides a comparative analysis of our proposed method against mainstream motion control techniques. In the first scenario involving the coordinated movement of two individuals, all methods manage to generate motion trajectories that are relatively accurate. However, our approach stands out for its superior visual quality. This advantage is largely attributed to the use of longer sequence frames, which contribute to smoother motion trajectories and more realistic background rendering. For example, in our generated bicycle scenario, the human legs exhibit lifelike pedaling motions, while the output from DragNUWA shows legs that appear to float almost horizontally, compromising physical realism. Moreover, both DragNUWA and MotionCtrl encounter significant motion blur towards the end of their videos. Additionally, MotionCtrl introduces unintended camera movements during the riding sequence, despite the absence of any intended camera movement conditions. In another instance, DragNUWA suffers from severe deformation of the lantern as the provided trajectory oscillates up and down. Although MotionCtrl’s trajectory is relatively accurate, the resulting video does not align with the expected portrayal of two lanterns. Overall, our method not only adheres closely to the provided trajectories but also minimizes object deformation, thereby ensuring higher fidelity in motion representation.
Compare with OpenSora
Despite OpenSora’s impressive accomplishments, it faces challenges when creating long videos featuring complex motions, such as simultaneous movement of multiple objects, swinging, or circling. This often leads to incoherent or distorted foreground objects, negatively impacting visual quality. To our delight, we discovered that incorporating appropriate trajectory control into the DiT model offers a more effective constraining signal. This improvement markedly enhances video fluidity and preserves object fidelity, as demonstrated in Figure  6.
@@
In scenarios where a teddy bear is oscillating side to side on a skateboard or a rose is swirling in circular motions, OpenSora, which relies solely on textual directives for motion control, exhibits noticeable object deformations. In contrast, Tora excels at maintaining the inherent shape of the objects. Additionally, when managing the motion of multiple entities, such as a pair of jellyfish—one moving upward while the other moves downward, OpenSora demonstrates noticeable flickering, underscoring its limitations in handling complex movements. In conclusion, the integration of Tora’s motion signaling mechanism enhances both the controllability and stability of the synthesized videos.
Comparison of Different Trajectory Compression Methods
@@
Data Pre-processing
@@
Initially, during the dataset preparation phase, we remove invalid videos. This step aims to identify and discard videos that do not meet our established criteria, including those with encoding errors, a duration of zero, or low quality. We identify encoding errors and zero-duration videos by directly decoding them. Furthermore, we predict both the aesthetic score [^4] https://github.com/christophschuhmann/improved-aesthetic-predictor and the optical flow score (Xu et al. 2023) for each video. A video is deemed valid only if its aesthetic score exceeds 5.5 and its flow score is greater than 2.
@@
Subsequently, we perform camera motion filtering using a camera motion detector [^5] https://github.com/antiboredom/camera-motion-detector and a motion segmentor  (Zhao et al. 2022) to filter out videos with significant camera movement, which may distort the model’s ability to focus on the motion of the primary subjects. More specifically, the zoom detection threshold is set between 0.4 and 0.6. The detected camera movement angles, calculated based on the background from the motion segmentation results, are valid as follows:$[0^{\circ},20^{\circ}],[160^{\circ},200^{\circ}],[340^{\circ},360^{\circ}]$.
Finally, we analyze the magnitude score of the optical flow within the foreground, excluding those scenes that are mostly static or exhibit minimal movement. Moreover, dramatic object motions in some videos can cause significant optical flow deviations, interfering with trajectory training. Consequently, we retain these videos with a probability of $(1-score/100)$.
@@
Evaluation Data
Our evaluation dataset is primarily sourced from video object segmentation datasets (Xu et al. 2021; Qi et al. 2022; Pont-Tuset et al. 2017), which offer robust object motion critical for our analysis. To enhance the quality of our evaluation, we implement a camera motion filtering technique to select videos where the camera remains predominantly stable. This filtering process allows us to concentrate on where object motion is distinctly pronounced, thereby improving the reliability of our assessments. For each frame, we utilize the center of the annotated object masks as trajectory points, providing precise references for evaluating motion dynamics. Figure 9 presents several examples from our evaluation dataset, highlighting the diversity and relevance of the selected video sequences.
Prompt Refinement
@@
Motion VAE Training
Given the absence of pre-existing networks tailored for video optical flow compression, training such a network from scratch presents significant challenges. Directly transferring the motion vectors to the image domain and applying a pretrained 3D VAE may hinder the model’s ability to effectively encode motion features, primarily due to domain discrepancies. To overcome this issue, we refine a motion-specific 3D VAE that is initialized from a pretrained model. Specifically, our motion 3D VAE is specifically initialized using the architecture of OpenSora’s VAE, which adapts the structure of Magvit-v2. This VAE has a substantial parameter count of 384 million, effectively leveraging the capabilities of a well-established network. Our training data is sourced from a combination of datasets annotated with optical flow information (Mehl et al. 2023; Mayer et al. 2016; Ranjan et al. 2020; Cabon, Murray, and Humenberger 2020). We fine-tune the motion 3D VAE for 200,000 iterations with a batch size of 1. The training video size is set to a random number of frames, capped at 34. This setting aligns with the OpenSora video VAE, improving compatibility between the motion VAE and the video VAE and ensuring a cohesive training process. We utilize PSNR, SSIM and trajectory error to evaluate reconstruction quality and motion-controllable ability. The performance differences between the pure video VAE and our fine-tuned model are presented in Table 7.

### v4

@@ 4.2 Results
UNet-based method
VideoComposer [39] 529 668 856 0.2335 0.2284 0.2236 15.11 29.14 58.76
DragNUWA [48] 475 593 784 0.2385 0.2341 0.2305 10.04 17.33 41.25
AnimateAnything [9] 487 602 775 0.2399 0.2342 0.2313 13.39 27.28 51.33
TrailBlazer [20] 459 581 756 0.2403 0.2351 0.2322 11.68 19.47 44.10
MotionCtrl [41] 463 572 731 0.2412 0.2376 0.2331 9.42 16.46 38.39
DiT-based method
OpenSora [55] 430 476 533 0.2452 0.2433 0.2411 286.43 321.52 373.17
OpenSora-based DragNUWA* 451 504 565 0.2430 0.2419 0.2393 10.11 13.88 21.75
@@
Table 1: Quantitative comparisons with motion-controllable video generation models. As the number of generated frames increases, Tora’s performance advantage over UNet-based methods becomes more pronounced. Specifically, Tora not only enhances motion fidelity but also improves the visual quality of the foundational model. Comparisons with OpenSora-based DragNUWA highlight the strengths of our proposed motion modules, which integrate seamlessly with DiT’s architecture.
We evaluate against motion-guided video generation methods using 16-frame, 64-frame, and 128-frame configurations. Trajectories are scaled proportionally for different durations. UNet-based approaches utilize sequential inference for extended generation. Since there are no DiT-based methods, we adapt DragNUWA’s motion trajectory design to our foundation model as an additional baseline. Specifically, we implement its official convolutional motion feature extraction and linear projection-based injection. To ensure compatibility with base DiT model, we perform downsampling across both spatial and temporal dimensions.
[…]
@@
Figure 4: Comparison of Trajectory Error across various resolutions and durations.
Figure 4 presents an analysis of Trajectory Error across different resolutions and durations. Unlike UNet models, Tora shows a gradual increase in error as duration extends. This aligns with the decrease in video quality observed in the DiT model. The results demonstrate that our method effectively maintains trajectory control over longer durations.
@@
Table 2: Evaluation of the impact of different trajectory compression methods.
Figure 5: Qualitative Comparisons on Trajectory Control. In the bicycle scenario, Tora realistically captures pedaling motions, while other methods show legs in an unnatural, nearly horizontal position. In another case, DragNUWA causes significant deformation of the lanterns, and MotionCtrl fails to accurately depict two lanterns. Overall, Tora not only adheres precisely to the specified trajectory but also produces smoother movement that conforms to the physical world.
4.3 Ablation study
Trajectory Compression. To validate the alignment of motion patches with the DiT inputs, we explore three different methods for trajectory compression, as summarized in Table 2. The first method samples the mid-frame as a keyframe for successive 4-frame intervals and uses Patch-Unshuffle [15] for spatial compression. While simple, this approach is sub-optimal for motion control due to potential flow estimation errors during rapid movements or occlusions, and the increased dissimilarity between patches complicates learning. The second method employs average pooling to gather information from successive frames. Although this captures a general sense of movement, it sacrifices precision by averaging the trajectory’s direction and magnitude, diluting important motion details. Our method utilizes a 3D motion VAE to extract the global context of successive trajectory intervals. Extensive training on a large dataset of trajectory videos with this method yields the best results, highlighting the effectiveness of our customized VAE approach for compression.
@@ 4.3 Ablation study
Block design and integrated position of MGF. We train the three variant MGF blocks as previously described, with the results presented in Table 3. Notably, the adaptive norm block achieves lowest FVD and Trajectory Error, while also exhibiting the highest computational efficiency. This advantage arises from its ability to dynamically adapt features based on varying conditions without strict alignment, a common cross-attention challenge. It also maintains temporal consistency by modulating conditional information over time, essential for incorporating motion cues. In contrast, channel concatenation can lead to information congestion, making motion signals less effective. We find that initializing the normalization layer as the identity function is vital for optimal performance. Additionally, placing the MGF module within the Temporal DiT block significantly enhances trajectory motion control, evidenced by a drop in Trajectory Error from 23.39 to 14.25.
@@
Table 4: Ablation of the type of training trajectories. “Hybrid" denotes the two-stage training strategy.
Training Strategies. We evaluate the two-stage training approach, with results in Table 4. Training only with dense optical flow is ineffective, as it fails to capture the details of sparse trajectories. Conversely, using only sparse trajectories provides limited information, complicating the learning process. In contrast, our two-stage strategy demonstrates better adaptability and versatility in managing various motion patterns, leading to improved overall performance.
Scaling motion-control ability. We investigate scaling laws in motion-controllable video generation by examining model parameter size and training volume. To leverage existing multi-size foundational models, we transfer our motion modules to CogVideoX architectures (2B/5B parameters), resulting in Tora-CogVideoX2B (2.5B) and Tora-CogVideoX5B (6.3B). Our implementation retains base VAE for motion compression while inserting MGF prior to Expert Transformer’s full attention module. Figure 6 demonstrates that increasing model scale and training data improves motion control well, confirming our modules’ seamless compatibility with DiT’s scalability framework.
Figure 6: Scaling behavior of motion control ability in Tora.
5 Conclusion
This paper introduces Tora, the first trajectory-oriented Diffusion Transformer framework for video generation. Tora effectively encodes arbitrary trajectories into spacetime motion patches, which align well with the scaling properties of DiT, thereby enabling more realistic simulations of physical world movements. By employing a two-stage training process, Tora achieves motion-controllable video generation across a wide range of durations, aspect ratios, and resolutions. Remarkably, it can generate high-quality videos that adhere to specified trajectories, producing up to 204 frames at 720p resolution. This capability underscores Tora’s versatility and robustness in handling diverse motion patterns while maintaining high visual fidelity. We hope our work provides a strong baseline for future research in motion-guided Diffusion Transformer methods.
@@ 5 Conclusion
6 Qualitative Comparisons
While the main text focuses on quantitative comparisons with the motion-controllable video generation models and ablation studies on different designs for Trajectory Extractor and Motion-guidance Fuser, here we provide further visual comparisons.
@@ 6 Qualitative Comparisons
6.1 Compare with OpenSora
Despite OpenSora’s impressive accomplishments, it faces challenges when creating long videos featuring complex motions, such as simultaneous movement of multiple objects, swinging, or circling. This often leads to incoherent or distorted foreground objects, negatively impacting visual quality. To our delight, we discovered that incorporating appropriate trajectory control into the DiT model offers a more effective constraining signal. This improvement markedly enhances video fluidity and preserves object fidelity, as demonstrated in Figure  7.
@@ 6.1 Compare with OpenSora
In scenarios where a teddy bear is oscillating side to side on a skateboard or a rose is swirling in circular motions, OpenSora, which relies solely on textual directives for motion control, exhibits noticeable object deformations. In contrast, Tora excels at maintaining the inherent shape of the objects. Additionally, when managing the motion of multiple entities, such as a pair of jellyfish—one moving upward while the other moves downward, OpenSora demonstrates noticeable flickering, underscoring its limitations in handling complex movements. In conclusion, the integration of Tora’s motion signaling mechanism enhances both the controllability and stability of the synthesized video output.
6.2 Comparison of Different Trajectory Compression Methods
@@
7 User Study with DiT-based methods
We conduct a user study to compare OpenSora-v1.2, CogVideoX-2B [47], Vidu [3], and Kling v1.0 [1], assessing the effectiveness of our method using our evaluation dataset. 10 human volunteers participate in evaluating quality based on three criteria: physics simulation, sensory quality, and instruction adherence. For Tora, participants draw appropriate trajectories in response to given text prompts. The experiment employs a pairwise comparison approach, where evaluators choose the superior output from each pair of generated results based on the same input. The resulting win rates appear in Table 5. Our method outperforms both OpenSora and CogVideoX-2B across all metrics, affirming the superiority of our proposed modules and data processing methods. Compared to the closed-source method, Vidu, we achieve competitive results. Kling demonstrates remarkable capabilities, and we hope that Tora can work to close the performance gap in future iterations.
Model Phys. Simu. Sens. Qual. Inst. Foll.
Tora vs. OpenSora-v1.2 71% 61% 64%
Tora vs. CogVideoX 53% 56% 52%
Tora vs. Vidu 54% 48% 47%
Tora vs. Kling 45% 43% 41%
Table 5: Win rates of Tora compared to OpenSora-v1.2, CogVideoX, Vidu, and Kling in terms of Physics Simulation, Sensory Quality, and Instruction Following.
8 Data Pre-processing
@@
Initially, during the dataset preparation phase, we remove invalid videos. This step aims to identify and discard videos that do not meet our established criteria, including those with encoding errors, a duration of zero, or low quality. We identify encoding errors and zero-duration videos by directly decoding them. Furthermore, we predict both the aesthetic score [^4] https://github.com/christophschuhmann/improved-aesthetic-predictor and the optical flow score [44] for each video. A video is deemed valid only if its aesthetic score exceeds 5.5 and its flow score is greater than 3.
@@
Subsequently, we perform camera motion filtering using a camera motion detector [^5] https://github.com/antiboredom/camera-motion-detector and a motion segmentor  [54] to filter out videos with significant camera movement, which may distort the model’s ability to focus on the motion of the primary subjects. More specifically, the zoom detection threshold is set between 0.4 and 0.6. The detected camera movement angles, calculated based on the background from the motion segmentation results, are valid as follows:$[0^{\circ},20^{\circ}],[160^{\circ},200^{\circ}],[340^{\circ},360^{\circ}]$.
Finally, we analyze the magnitude score of the optical flow within the foreground, excluding those scenes that are mostly static or exhibit minimal movement. Moreover, dramatic object motions in some videos can cause significant optical flow deviations, interfering with trajectory training. Consequently, we retain these videos with a probability of $(1-flow\_score/100)$.
@@ 9.1 Training Data
9.2 Evaluation Data
Our evaluation dataset is primarily sourced from video object segmentation datasets [46, 31, 30], which offer robust object motion critical for our analysis. To enhance the quality of our evaluation, we implement a camera motion filtering technique to select videos where the camera remains predominantly stable. This filtering process allows us to concentrate on where object motion is distinctly pronounced, thereby improving the reliability of our assessments. For each frame, we utilize the center of the annotated object masks as trajectory points, providing precise references for evaluating motion dynamics. Figure 10 presents several examples from our evaluation dataset, highlighting the diversity and relevance of the selected video sequences.
10 Prompt Refinement
@@
11 Motion VAE Training
Given the absence of pre-existing networks tailored for video optical flow compression, training such a network from scratch presents significant challenges. Directly transferring the motion vectors to the image domain and applying a pretrained 3D VAE may hinder the model’s ability to effectively encode motion features, primarily due to domain discrepancies. To overcome this issue, we refine a motion-specific 3D VAE that is initialized from a pretrained model. Specifically, our motion 3D VAE is specifically initialized using the architecture of OpenSora’s VAE, which adapts the structure of Magvit-v2. This VAE has a substantial parameter count of 384 million, effectively leveraging the capabilities of a well-established network. Our training data is sourced from a combination of datasets annotated with optical flow information [22, 21, 33, 6]. We fine-tune the motion 3D VAE for 200,000 iterations with a batch size of 1. The training video size is set to a random number of frames, capped at 34. This setting aligns with the OpenSora video VAE, improving compatibility between the motion VAE and the video VAE and ensuring a cohesive training process. We utilize PSNR, SSIM and Trajectory Error to evaluate reconstruction quality and motion-controllable ability. The performance differences between the pure video VAE and our fine-tuned model are presented in Table 8.